# Understanding the Limits of Lifelong Knowledge Editing in LLMs

### v1

@@ 4 Understanding the Limits of Lifelong Knowledge Editing in LLMs
Experimental Details. All experiments are performed on a compute cluster equipped with Nvidia A100 and H100 GPUs leveraging PyTorch (Paszke et al., 2019) and building on the EasyEdit codebase (Zhang et al., 2024). All our evaluations are conducted on five state-of-the-art LLMs: Llama-2-7b (Touvron et al., 2023b) Llama-3-8b (Grattafiori et al., 2024), Mistral-7b (Jiang et al., 2023), Gemma-7b (Team et al., 2024a) and xGen-7b (Nijkamp et al., 2023). To representatively study knowledge editing, we leverage the following approaches: ROME (Meng et al., 2023), R-ROME (Gupta et al., 2024a), MEMIT (Meng et al., 2022) and WISE (Wang et al., 2024a). For our retrieval-augmented generation experiments, we store update question-answer pairs in memory alongside associated key embeddings generated by embedding questions with the all-mpnet-base-v2 embedding model (Song et al., 2020), the best general sentence-embedding model provided in the SBERT repository (Reimers & Gurevych, 2019). During inference, the top-2 most similar entries are retrieved from memory efficiently using the Annoy solver (Bernhardsson, 2018). This setup ensures efficient retrieval and supports multi-hop questions with minimal computational overhead. For our continual finetuning experiments, we explore variations of low-rank adaptation using LoRA (Hu et al., 2021) alongside simple interpolation-based model-merging (Wortsman et al., 2022; Stojanovski et al., 2022; Marouf et al., 2024) which has shown to perform competitively even when compared to more involved merging techniques (Roth et al., 2024; Dziadzio et al., 2024). Adapters are trained for 10 epochs on each timestep. Additional details are provided in the corresponding subsections and the supplementary.
@@ 4.1 Lifelong Knowledge Editing on WikiBigEdit
To begin, Figure 3 presents a high-level comparison of the average accuracy across all eight timesteps $[U_{1},...,U_{8}]$ for all five LLMs listed above. After incorporating factual changes of a timestep, evaluation is conducted across all five evaluation axes spanned by respective evaluation data $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$, $\mathcal{S}^{\text{QA}}_{\text{personas}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$, and a performance evaluation on the actual edit data $\mathcal{S}^{\text{QA}}_{\text{update}}$. Results across all models and timesteps are averaged and visualized.
@@
RAG naturally trades performance increases for longer inference times. However, with effective approximate solvers and high-quality sentence embeddings, we find that inference time at most doubles after processing all of WikiBigEdit (c.f. Figure 16). Even so, knowledge editing techniques like WISE and MEMIT, specifically designed to factually update LLMs, cannot match simple low-rank adapter-based finetuning (LoRA-FT) at scale. Moreover, with interpolation-based model merging (LoRA-Merge), commonly used in training and finetuning LLMs (Yang et al., 2024a; Roth et al., 2024; Team et al., 2024a, b; Dziadzio et al., 2024), consistent gains are observed across all evaluation axes. These results strongly suggest that at scale, standard knowledge editing techniques fall short even when compared to straightforward continual finetuning approaches commonly employed for general continual pretraining (Ibrahim et al., 2024; Roth et al., 2024).
@@
The following sections provide a more detailed breakdown into specific failure scenarios and drivers.
@@ 4.2 The Limits of Knowledge Editing at Scale
Key Takeaway Knowledge editing reliant on local weight modification collapses quickly. Even methods accounting for lifelong edits simply converge to pre-update performance early. Current editing methods appear unsuitable for large-scale sequential knowledge editing.
@@ 4.3 Retrieval Augmentation for WikiBigEdit
Looking at the detailed results of the RAG baseline on WikiBigEdit in Figure 5, the improvements across all evaluation axes (except multi-hop reasoning) are significant across 500K+ updates. The top row shows the average accuracy on the current timestep, with bold lines representing average performance across five models (light background lines). The bottom row illustrates average performance across all timesteps after training on the current timestep’s edits. As noted, RAG achieves near-perfect accuracy on $\mathcal{S}^{\text{QA}}_{\text{update}}$ by retrieving stored questions from memory, showing no signs of “forgetting” on trained edits (c.f. bottom row, leftmost). Additionally, edits remain highly localized since no weight updates are performed. While RAG maintains strong performance on generalization axes (“rephrasing” at 78.64% and “personas” at 66.16%), it still exhibits symptoms of catastrophic forgetting. Performance on older timestep data declines with each new large update batch (bottom, second and third), as retrieval memory becomes increasingly obfuscated by irrelevant context for specific timesteps.
Finally, while RAG outperforms knowledge editing approaches on standard measures (locality, generalization, and overall edit performance), its multi-hop reasoning accuracy remains near pre-update levels (second from right in Figure 5). Figure 6 reveals the bottleneck: retrieval accuracy is low - likely due to the absence of the middle subject in the multi-hop question - with correct retrieval rates below 50% for the first hop and even lower for the second. Only 10% of evaluation samples retrieve both facts correctly. Even when both hops are retrieved, answer accuracy remains near single-hop levels (i.e. solving multi-hop questions using only information from just the first hop), as without training, models struggle to combine retrieved information without training for reasoning tasks.
@@ 4.4 Continual Finetuning
Interestingly, coupling continual LoRA finetuning with simple model merging (interpolating adapter weights with a 0.25 weight after each timestep) significantly improves editing performance across WikiBigEdit. While LoRA-Merge initially lags on shorter update sequences, it maintains superior stability as updates accumulate, surpassing continual finetuning - and thereby all specialized knowledge editing techniques - after approximately 100K edits. Notably, LoRA-Merge exhibits minimal catastrophic forgetting and even shows signs of positive backward transfer, where training on new updates improves retention of prior edits. These findings reinforce the effectiveness of model merging in continual finetuning and pretraining (Stojanovski et al., 2022; Thede et al., 2024; Roth et al., 2024), highlighting its promise for large-scale knowledge editing.
@@ 5 Conclusion
Leveraging WikiBigEdit, we conduct an extensive experimental study in which we identify key limitations of existing knowledge editing approaches such as ROME, MEMIT, and WISE, which struggle with scalability and stability in large-scale update settings. Our results also demonstrate the effectiveness of retrieval-augmented generation (RAG) when no restriction on inference time is given and highlight continual finetuning with model merging as a competitive alternative at an equivalent inference cost.
@@ A.3 Persona Set
Table 3: Descriptions of the personas used for rephrasing questions with varied styles.

### v2

@@ 4 Understanding the Limits of Lifelong Knowledge Editing in LLMs
Experimental Details. All experiments are performed on a compute cluster equipped with Nvidia A100 and H100 GPUs, leveraging PyTorch (Paszke et al., 2019) and building on the EasyEdit codebase (Zhang et al., 2024). All our evaluations are conducted on five state-of-the-art LLMs: Llama-2-7b (et al., 2023) Llama-3-8b (et al., 2024a), Mistral-7b (Jiang et al., 2023), Gemma-7b (Team et al., 2024a) and xGen-7b (Nijkamp et al., 2023). To representatively study knowledge editing, we leverage the following approaches: ROME (Meng et al., 2023), R-ROME (Gupta et al., 2024a), MEMIT (Meng et al., 2022), and WISE (Wang et al., 2024a). For our retrieval-augmented generation experiments, we store update question-answer pairs in memory alongside associated key embeddings generated by embedding questions with the all-mpnet-base-v2 embedding model (Song et al., 2020), the best general sentence-embedding model provided in the SBERT repository (Reimers & Gurevych, 2019). During inference, the top-2 most similar entries are retrieved from memory efficiently using the Annoy solver (Bernhardsson, 2018). This setup ensures efficient retrieval and supports multi-hop questions with minimal computational overhead. For our continual finetuning experiments, we explore variations of low-rank adaptation using LoRA (Hu et al., 2021) alongside simple interpolation-based model-merging (Wortsman et al., 2022; Stojanovski et al., 2022; Marouf et al., 2024) which has shown to perform competitively even when compared to more involved merging techniques (Roth et al., 2024; Dziadzio et al., 2024). Adapters are trained for 10 epochs on each timestep. Additional details are provided in the corresponding subsections and the supplementary.
@@ 4.1 Lifelong Knowledge Editing on WikiBigEdit
To begin, Figure 3 presents a high-level comparison of the average accuracy across all eight timesteps $U_{1},...,U_{8}]$ for all five LLMs listed above. After incorporating factual changes of a timestep, evaluation is conducted across all five evaluation axes spanned by respective evaluation data $\mathcal{S}^{\text{QA}}_{\text{rephrased}}$, $\mathcal{S}^{\text{QA}}_{\text{personas}}$, $\mathcal{S}^{\text{QA}}_{\text{locality}}$, $\mathcal{S}^{\text{QA}}_{\text{mhop}}$, and a performance evaluation on the actual edit data $\mathcal{S}^{\text{QA}}_{\text{update}}$. Results across all models and timesteps are averaged and visualized.
@@
RAG naturally trades performance increases for longer inference times. However, with effective approximate solvers and high-quality sentence embeddings, we find that inference time at most doubles after processing all of WikiBigEdit (c.f. LABEL:fig:forward_pass_time). Even so, knowledge editing techniques like WISE and MEMIT, specifically designed to factually update LLMs, cannot match simple low-rank adapter-based finetuning (LoRA-FT) at scale. Moreover, with interpolation-based model merging (LoRA-Merge), commonly used in training and finetuning LLMs (Yang et al., 2024a; Roth et al., 2024; Team et al., 2024a, b; Dziadzio et al., 2024), consistent gains are observed across all evaluation axes. These results strongly suggest that at scale, standard knowledge editing techniques fall short even when compared to straightforward continual finetuning approaches commonly employed for general continual pretraining (Ibrahim et al., 2024; Roth et al., 2024).
@@
The following sections provide a more detailed breakdown of specific failure scenarios and drivers.
@@ 4.2 The Limits of Knowledge Editing at Scale
Key Takeaway Knowledge editing is reliant on local weight modification and collapses quickly. Even methods designed for lifelong editing converge to pre-update performance early. So large-scale lifelong editing requires significantly improving existing methods.
@@ 4.3 Retrieval Augmentation for WikiBigEdit
Looking at the detailed results of the RAG baseline on WikiBigEdit in Figure 5, the improvements across all evaluation axes (except multi-hop reasoning) are significant across 500K+ updates. The top row shows the average accuracy on the current timestep, with bold lines representing average performance across five models (light background lines). The bottom row illustrates average performance across all timesteps after training on the current timestep’s edits. As noted previously, RAG achieved near-perfect accuracy on $\mathcal{S}^{\text{QA}}_{\text{update}}$ by simply retrieving the exact stored questions from memory; exhibiting no sign of “forgetting” on trained edits (c.f. bottom row, leftmost). Additionally, edits remain highly localized since no weight updates are performed. Interestingly, while RAG also maintains high average performance on generalization-based evaluation axes (“rephrasing” and “personas”), achieving average accuracies of 78.64% and 66.16%, respectively, RAG still exhibits symptoms akin to catastrophic forgetting encountered in continual learning and knowledge editing (Kirkpatrick et al., 2017; Zenke et al., 2017; Wang et al., 2024a). Here, performance on older timestep data decays with every new large update batch (bottom, second and third), as the retrieval memory becomes more and more obfuscated with context irrelevant for data from particular timesteps.
Finally, while RAG outperforms knowledge editing approaches on standard measures (locality, generalization, and overall edit performance), its multi-hop reasoning accuracy remains near pre-update levels (second from right in Figure 5). Figure 6 reveals the bottleneck: retrieval accuracy is low - likely due to the absence of the middle subject in the multi-hop question - with correct retrieval rates below 50% for the first hop and even lower for the second. Only 10% of evaluation samples retrieve both facts correctly. Even when both hops are retrieved, answer accuracy remains near single-hop levels (i.e. solving multi-hop questions using only information from just the first hop), as models struggle to combine retrieved information without dedicated training for reasoning tasks. While advanced methods like HippoRAG (Gutiérrez et al., 2025) may help, they typically require longer contexts or higher compute; we leave their integration to future work.
@@ 4.4 Continual Finetuning
Interestingly however, as we couple continual LoRA finetuning with simple model merging (after training of each timestep, current adapter weights are simply merged into preceding adapter weights using an interpolation weight of 0.25), we achieve significantly improved editing performance across WikiBigEdit. While LoRA-Merge initially shows lower performance on shorter update sequences, it demonstrates superior stability as updates accumulate; and begins to notably outperform simple continual finetuning (and thereby also any specialized knowledge editing technique) after approximately 100K edits on all evaluation sets, maintaining consistent performance without significant degradation. Importantly, LoRA-Merge shows at most marginal evidence of catastrophic forgetting, and in parts even exhibits positive backward transfer (where training on a current timestep actually improves performance on all edits seen previously) - see Figure 7 bottom row. These results reinforce the effectiveness of model merging for general continual finetuning and pretraining tasks (Stojanovski et al., 2022; Thede et al., 2024; Roth et al., 2024), highlighting its promise for large-scale knowledge editing.
@@ 5 Conclusion
Leveraging WikiBigEdit, we conduct an extensive experimental study in which we identify key limitations of existing knowledge editing approaches such as ROME, MEMIT, and WISE, which struggle with scalability and stability in large-scale update settings. Our results also demonstrate the efficacy of retrieval-augmented generation (RAG) when no restriction on inference time is given and highlight continual finetuning coupled with model merging as a competitive alternative at an equivalent inference cost.
@@ A.3 Persona Set
Table 4: Results of WISE on three language models (Llama-2-7B, Llama-3-8B, and Mistral-7B) for the first timestep of the WikiBigEdit benchmark. Each subplot shows mean batch accuracy for the update, rephrase, personas, mhop, and locality sets over the number of updates. Post-update performance (blue) converges to pre-update levels (gray) across all metrics, highlighting WISE’s limitations in sustaining accuracy after large-scale updates.